OpenStudy (anonymous):

if y=2cos(x/2), then (d^2y)/(dx^2)=

pooja195 (pooja195):

y = 2cos(x/2) y' = -2(1/2) sin (x/2) = -sin(x/2) y'' = -(1/2) cos (x/2)

OpenStudy (anonymous):

-sin(x/2)

OpenStudy (anonymous):

ok we got 2 different answers here

OpenStudy (anonymous):

-(1/2)cos(x/2)

OpenStudy (anonymous):

there we go thank you